Suppose there is an investment proposal requiring $16,000 outlay
now (time zero) and returning a constant cash flow of $7,000 per
period before tax savings due to interest payments for the next
three years. The proposal is to have a market debt proportion of
50% (i.e., 0.50). The capital market requires per period rate of
return on equity of 27% and on debt of 9%. The corporate tax rate
is 40%, and interest is deductible for the calculation of income tax.
Calculate the NPV and IRR based on (a) weighted average cost
of capital method, (b) Arditty-Levy method, (c) equity residual
method, and (d) adjusted net present value method.
1
Expert's answer
2019-08-29T09:22:19-0400

(a) WACC is calculated by multiplying the cost of each capital source (debt and equity) by its relevant weight, and then adding the products together to determine the value.

WACC = 0.27×0.5 + 0.09×0.5 = 0.18.

So, NPV = -16,000 + 7,000×(1 - 0.4)/1.18 + 7,000×(1 - 0.4)/1.18^2 + 7,000×(1 - 0.4)/1.18^3 = -6868.05.

IRR is the discount rate at which NPV = 0.

In our case NPV < 0 even if IRR = 0, and IRR can't be negative, so it is impossible to find IRR in this case.

(b) Arditty-Levy method is often used by public utilities. It uses EBIT, corporate tax rate, and depreciation.

(c) Equity residual method is used in banking and for international investments.

(d) Adjusted present value (APV), defined as the net present value of a project if financed solely by equity plus the present value of financing benefits, is another method for evaluating investments. It is very similar to NPV. The difference is that is uses the cost of equity as the discount rate rather than WACC.